In the style of the KA characterization of the CM board, the AO path gain EPICS slider (IN2) of the IMC servo board was stepped by 1 dB through the full available range of -32 dB to +31 dB. For each value of the requested gain, I measured the TF from the injected signal (to IN2) to TP1A on the IMC servo board. I used the BNC connector for this test, whereas we use the LEMO connector for the AO path. The source was tee-d off at the SR785 side, with one leg going to IN2 of the IMC servo board, and the other going to CH1A of the SR785. TP1A of the IMC board was connected to CH2A of the SR785.
Attachment #1 - Measured gain vs requested gain.
Attachment #2 - Frequency dependent transfer functions
The motivation here is to try and figure out why I cannot engage the AO path smoothly in the CARM handoff part of lock acquisiton. I plan to use this information to do some loop modeling and project laser frequency noise coupling in various stages of the lock acquisition process.